OpenStudy (anonymous):

. A mini license plate for a toy car must consist of three numbers followed by a letter. Each number must be a 3, 6, or 9. Repetition of digits is NOT permitted. Each letter must be an A, B or C. Use the counting principle to determine the number of points in the sample space. List the sample space. Determine the exact probability of creating a mini license plate with a B. Give solution exactly in reduced fraction form.

OpenStudy (anonymous):

= 3 x2 x 1 x 3 = 18ways

OpenStudy (anonymous):

probability of having a B = 6/18 = 1/3